Q.1Alexander Company purchased a piece of equipment for $12,000 and depreciated it for three years over a five-year estimated life with an expected residual value at the end of five years of $2,000. At the end of the third year, Alex decided to upgrade to equipment with increased capacity and sold the original piece of equipment for $7,200. Calculate the gain or loss on the disposal at the end of the third year.



1
Expert's answer
2021-01-08T11:20:44-0500

Answer:

The value of an asset decreases with usage and time. This decreased value is called depreciation.


Cost of equipment is $12,000

Estimated life is 5 years

Residual value is $2000

"Depreciation = (cost of equipment - residual life)\/estimated life"

            = (12000 - 2000) \ 5

            = $2000 per year


Value of equipment at the end of first year is $12000 - 2000 = $10000

Value of equipment at the end of 2nd year is $10000 - 2000 = $8000

Value of equipment at the end of 3rd year is $8000 - 2000 = $6000


Sale price of equipment at the end of 3rd year = $7,200

Book value of equipment at the end of 3rd year = $6000

Gain on disposal of equipment = 7200 - 6000 

                        = $1200


Therefore, the gain on disposal of equipment is $1200